Write and equation and solve the following problem.
The area or a rectangle is 130 square feet. Find the dimensions of the rectangle if the length is 3 feet greater than the width.
I know A=LW so: 130 = L+3 = w
Did I get the formula correct

Answer by venugopalramana(3286)

SEE MY COMMENTS BELOW
The area or a rectangle is 130 square feet. Find the dimensions of the rectangle if the length is 3 feet greater than the width.
I know A=LW....OK
so: 130 = L+3 = w....NO....GO STEP BY STEP....LENGTH IS MORE THAN WIDTH...SO
L=W+3...OK...NOW AREA =LW=(W+3)W=130
SO WE GET
W^2+3W=130
W^3+3W-130=0
W^2+13W-10W-130=0
W(W+13)-10(W+13)=0
(W-10)(W+13)=0
W-10=0....OR...W=10
SO L=W+3=10+3=13
Did I get the formula correct